# Mental-Perceiver: Audio-Textual Multimodal Learning for Mental Health Assessment

### v1

@@ Conclusion
In this work, we construct a new large-scale Multi-Modal Psychological assessment corpus (MMPsy) about anxiety and depression in adolescents who use Mandarin. The MMPsy contains audios and extracted transcripts of responses from anxious/depressed and non-anxious/non-depressed adolescent volunteers. There are 7,758 cleaned interview data for anxiety detection and 4,266 for depression detection. We further propose a novel mental disorder estimation network, named Mental-Perceiver, to detect anxious/depressive mental states automatically according to users’ audio and corresponding transcripts. Extensive experiments on MMPsy and the public DAIC-WOZ show the effectiveness and superiority of our proposed Mental-Perceiver.

### v2

@@ Conclusion
In this work, we construct a new large-scale Multi-Modal Psychological assessment corpus (MMPsy) about anxiety and depression in adolescents who speak Mandarin. The MMPsy contains audios and extracted transcripts of responses from anxious/depressed and non-anxious/non-depressed adolescent volunteers. There are 7,736 cleaned interview data for anxiety detection and 4,247 for depression detection. We further propose a novel mental disorder estimation network, named Mental-Perceiver, to detect anxious/depressive mental states automatically according to users’ audio and corresponding transcripts. Extensive experiments on MMPsy and the public DAIC-WOZ show the effectiveness and superiority of our proposed Mental-Perceiver.
Acknowledgements
This work was supported in part by the National Natural Science Foundation of China (NSFC) under Grant No. 62206314, GuangDong Basic and Applied Basic Research Foundation under Grant No. 2022A1515011835, Science and Technology Projects in Guangzhou under Grant No. 2024A04J4388 and China Postdoctoral Science Foundation funded project under Grant No. 2021M703687.